OpenStudy (anonymous):

The force of attraction between two oppositely charged pith balls is 5.3 × 10^-6 newtons. If the charge on the two objects is 6.7 × 10^-9 coulombs, what is the distance between the two charges? (k = 9.0 × 10^9 newton·meter^2/coulombs^2) Select one of the options below as your answer: A. 3.0 × 10^-1 meters B. 7. 0 × 10^-1meters C. 6.7 × 10^-1 meters D. 8.2 × 10^-1 meters E. 9.6 × 10^-1 meters

OpenStudy (anonymous):

f=kQq/r*r

OpenStudy (anonymous):

how should I put it in for the kQq?

OpenStudy (anonymous):

k= 9 x 10^9 q=6.7 x 10^-9=Q

OpenStudy (anonymous):

F=9*10^9*10^-9*6.7

OpenStudy (anonymous):

idk what im doing i just got 60.3...?

OpenStudy (anonymous):

do i divide it by something?

OpenStudy (anonymous):

F=9*10^9*(10^-9*6.7)^2/r*r

OpenStudy (anonymous):

its a part

OpenStudy (anonymous):

@theEric

OpenStudy (anonymous):

i edited the problem try it out now

OpenStudy (theeric):

Are pith balls treated as point charges?

OpenStudy (anonymous):

columbs law..?

OpenStudy (souvik):

if the distance between the balls is >> the radius of ball... then the balls are treated as point charges

OpenStudy (anonymous):

im about to guess i cant do this

OpenStudy (anonymous):

I think all this data is wrong

OpenStudy (theeric):

I did some research and I see what the deal with the pith balls is. And I agree with @souvik . Your class might assume them to be point charges.

OpenStudy (souvik):

i got A as the answer

OpenStudy (anonymous):

eric do u agree?

OpenStudy (souvik):

i used coulomb's law

OpenStudy (theeric):

I got close to souvik's answer: 0.27609473353686493820576562279187m. @souvik , me too! Doing algebra in my head, and using a calculator.

OpenStudy (souvik):

hah Ha... you put all the digits here...:)@theEric

OpenStudy (anonymous):

im going with A thankyou

OpenStudy (theeric):

But is the force of attraction negative, in the question?

OpenStudy (theeric):

I don't know if Coulomb's law is the only thing that matters.

OpenStudy (anonymous):

thankyou all so much I got a 79% on the test <3

OpenStudy (anonymous):

I passed :)

OpenStudy (souvik):

you cheated !!!

OpenStudy (anonymous):

yes and no

OpenStudy (souvik):

you would try it yourself

OpenStudy (anonymous):

im just getting credit back for scchool cause I got a 58% 2nd semester

OpenStudy (anonymous):

I have an ap student sitting next to me he was having trouble on alot of my questions too

OpenStudy (theeric):

Yes, only if you stole the answers. No, if you did everything yourself. It is apparent that you took a test that allowed you internet freedom and time. Hopefully it did not forbid learning during the test (like and open-book test). If it did... shame on you... Otherwise, you did not cheat if you used only study tools, and not direct answers. I understand that the situation probably motivated you to use others' answers, but I hope you did not. The test should reflect your ability at the time of taking the test. As always, I really hope that you did not cheat, and I hope I didn't give too much away.

OpenStudy (theeric):

What I do, when I have to learn for a question, is ask to understand the material. If possible, I'd work a related question! Good luck! I'm glad you finished! Now you can relax!

OpenStudy (anonymous):

dude no I didnt really cheat I worked problems out myself if I was struggling and cant figure it out ill ahve people suggest answers desnt mean I just go a click that answer I make sure it right

OpenStudy (anonymous):

and thanks

OpenStudy (theeric):

There you go! If it's your knowledge, then that's not cheating. I think us putting direct input into the question is cheating a little, but you're right - you do the evaluating, and the learning. Still, I hope you could get the same grade or better! But I know how the school system is, and I know it's hard to learn about something you want to avoid! That's a pitfall of the system.

hero (hero):

Put it this way, If I tried something like this at my university, I'd be expelled on the spot.

hero (hero):

Keep in mind, this isn't the only question he (or she) posted.

OpenStudy (theeric):

That might be fair. Some schools have official guidelines laid out to define academic dishonesty. Professor's can lay out rules as well, as allowed by the guidelines if there are any. At any rate, I hope everything is okay.

OpenStudy (theeric):

This is all definitely something to consider the next time!